OpenStudy (anonymous):

Look at the formula for the area of a circle below. A = 3.14(pi)r2 Which of the following equations can be used to solve for the radius (r)?

OpenStudy (anonymous):

\[A=\pi r2\]

OpenStudy (anonymous):

orrr \[\pi r ^{2}\]

OpenStudy (sasogeek):

Well area of a circle is \(\pi r^2 \) so you can use that to find r if the area has been given

OpenStudy (anonymous):

What equation can be used to solve for the radius (r)?

OpenStudy (anonymous):

C = 2 Pi r divide both sides by 2*Pi C/(2 * Pi) = r

OpenStudy (sasogeek):

\[\huge A=\pi r^2 \ \ \text{which is the area of a cirlce} \]\[\huge C=2\pi r \ \text{circumference of a circle} \]

OpenStudy (sasogeek):

you basically have to make r the subject :)

OpenStudy (anonymous):

So R=?

OpenStudy (sasogeek):

if area is given, then \[\huge r=\sqrt{\frac{A}{\pi}} \] and when you have the circumference given, \[\huge r=\frac{C}{2\pi} \]

OpenStudy (anonymous):

ohhhh i see

OpenStudy (sasogeek):

:)